A cuboid is of dimensions 60 cm×54 cm×30 cm. How many small cubes with side 6 cm can be placed in the given cuboid?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Solution: Given, Length of cuboid, l = 60 cm, Breadth of cuboid, b = 54 cm and
Height of cuboid, h = 30 cm
We know that, Volume of cuboid = lbh = 60 ×54×30 = 97200 cm^3
And Volume of cube = (Side)^3
= 6×6×6 = 216 cm^3
Also,Number of small cubes = volume of cuboid / volume of cube
= 97200/216
= 450
Hence , required cubes are 450.

Thanks